OpenStudy (anonymous):

a painting measuring 10 in by 16 in is surrounded by a frame of uniform width. If the combined area of the painting and frame is 280 square inches dertermine the width of the frame

OpenStudy (anonymous):

assume the width is x. (10 + 2x)(16 + 2x) = 280 4x^2 + 52x + 160 = 280 4x^2 + 52x - 120 = 0 x^2 + 13x - 30 = 0 (x + 15)(x - 2) = 0 x = 2 (negative width makes no sense)

OpenStudy (anonymous):

thanks